# Sustainable Visions: Unsupervised Machine Learning Insights on Global Development Goals

### v1

@@ Abstract
The United Nations’ 2030 Agenda for Sustainable Development outlines 17 goals to address global challenges. However, progress has been slower than expected and, consequently, there is a need to investigate the reasons behind this fact. In this study, we used a novel data-driven methodology to analyze data from 107 countries (2000–2022) using unsupervised machine learning techniques. Our analysis reveals strong positive and negative correlations between certain SDGs. The findings show that progress toward the SDGs is heavily influenced by geographical, cultural and socioeconomic factors, with no country on track to achieve all goals by 2030. This highlights the need for a region specific, systemic approach to sustainable development that acknowledges the complex interdependencies of the goals and the diverse capacities of nations. Our approach provides a robust framework for developing efficient and data-informed strategies, to promote cooperative and targeted initiatives for sustainable progress.
@@ 1 Introduction
In 2015, the 2030 Agenda for Sustainable Development Goals (SDG) of the United Nations [1] was established for countries to follow 17 SDGs, 169 targets and 231 globally comparable indicators in their development process (for details, see Appendix A). Since then, the world has faced several set-backs, such as the COVID-19 pandemic, acceleration of climate change, and serious armed conflicts that have slowed down this process. for these reasons and due to the SDGs being variably interlinked or interrelated with each other, it seems difficult or even unlikely for the countries to achieve the SDGs by 2030. Moreover, in 2023 an independent group of scientists appointed by the UN to assess the progress of the world in achieving the SDGs and to recommend how to move forward, stated [1]: “The world is not on track to achieve any of the 17 SDGs and cannot rely on change to happen organically. Without accelerated action, the ambitious plan that the world signed up to in 2015 will fail.”
@@
Therefore, it seems obvious that we lack understanding of the complex inter-linkages between different SDGs and what are the drivers leading to their changes. This in turn points out the need for more research, which is made possible by the vast amount of data that the 231 global indicators of the 17 SDGs have produced so far. A thorough analysis of these data is likely to give us some insight into the correlations between, e.g., armed conflicts, climate change, social inequalities, global crime, etc., and would serve as a necessary precursor for the next steps towards sustainability.
@@
As direct interpretation of raw data and because the roles and interlinkages between SDGs are difficult to decipher, our aim in this study is to get a deeper understanding and insight into them using a modern data analysis approach. We will do this by using unsupervised machine learning methods as they can reveal groups of countries with similar behavior regarding the 17 SDG’s scores. For example, in the case of diversity, even subtle differences between countries otherwise quite similar should be considered when proposing a new policy agenda. All of this should provide valuable information to decision makers at the national and local levels to make well-informed policy choices to achieve the goals and align with the 2030 Agenda. The best practices for decision making involve identifying challenges, using data to form opinions, and employing risk analysis while taking into account environmental constraints and local conditions [2].
@@
Sustainability, like many other human endeavors, can be represented as a dynamic and complex network with non-linear interactions and interdependencies between the Sustainable Development Goals [3]. This makes it difficult to consider individual SDGs in isolation, as they are interconnected and influenced by the intricate web of environmental, political, economic, cultural and social factors. Recent progress in complex systems research has enhanced our ability to analyze sustainability using new systemic approaches [4]. Here we will do this by employing the network science approach and its tools while recognizing that cultural factors play an important role and are difficult to isolate within the global indicators.
@@
There have been studies concerning the interrelations amongst the SDGs [5, 6, 7, 8, 9]. One of our objectives is to use several techniques to study in detail these interlinkages, which can be positive or negative, indicating that progress on one goal can enhance or hamper progress on other goals. For a country, identifying these synergies and trade-offs before setting its policy priorities is crucial to avoid creating new problems while solving existing ones. This interesting issue of SDG interlinkings has been examined in a recent review studies [10], which concluded that our understanding of sustainability is still quite rudimentary.
@@
In our study, we will use an unsupervised machine learning multimethod approach with three objectives in mind: i) to examine the changes over time in the 17 SDGs for all countries, identifying possible classifications and discovering similarities and differences in their patterns; ii) to understand how the SDGs correlate across different countries; and iii) to investigate the collective dynamics of similar countries to measure the rate of progress toward the achievement of the SDGs has been generated in the 21stcentury. Our methods uncover distinct SDG patterns associated with well-defined clusters of countries, each showing different dynamical behaviors toward ideal scores.
@@ 2 Materials and methods
As the main data analysis method we choose unsupervised unsupervised machine learning techniques to explore and classify the dynamic patterns of each countries’ 17 SDGs from 2000 to 2022. This included three main steps: i) We first applied dimensionality reduction techniques, beginning with a linear projection, Principal Component Analysis (PCA) [11, 12], to detect global patterns, followed by t-Distributed Stochastic Neighbor Embedding (t-SNE) [13, 14], a graph-based algorithm, to project the multidimensional SDG indicators into a lower-dimensional space and identify local patterns; ii) We then utilized the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [15] to cluster similar countries based on the reduced data representation; iii) Finally, we estimated the characteristic dynamics of each cluster by evaluating the statistical distribution of Euclidean distances from the ideal SDG achievements and computing the mean distance for each cluster annually.
@@ 2.1 Data
The data in this study was collected from the “Online database for the Sustainable Development Report 2023”[16], including only countries with no missing data. The original data include 166 countries, but 15 lack information about Goal 1 (no poverty), 17 had no information about Goal 10 (reduce inequalities) and 40 countries that lack coastal regions do not have scores for Goal 14 (life below water). Therefore, only 107 countries were included in our analysis.
2.2 Principal Component Analysis (PCA)
@@
PCA was performed using standardized data, i.e., means and standard deviations were calculated for each SDG across the years 2000 to 2022. The corresponding mean was then subtracted from each observation, and the result was divided by the SDG’s standard deviation.
2.3 t-Distributed Stochastic Neighbor Embedding (t-SNE)
The t-SNE [13, 14] algorithm is a powerful technique for visualizing high-dimensional data by mapping it into a lower-dimensional space, typically 2 or 3 dimensions. The primary objective of t-SNE is to represent the underlying structure of the data by placing similar data points close together and dissimilar points further apart in the lower-dimensional space. This is achieved by converting the pairwise Euclidean distances between data points into conditional probabilities, which reflect the likelihood of points being neighbors in the original high-dimensional space. The resulting visualization preserves the local relationships within the data, making it easier to identify patterns and clusters. A more in depth explanation of the method can be found in Appendix  D.
@@ 3.1 Global view of SDG scores
Given that the effectiveness of sustainability relies on global collaboration, our initial findings focus on the average Sustainable Development Goals (SDGs) scores at a global scale. Figure  1 illustrates the yearly mean scores for each goal, enabling simultaneous observation of the values of all 17 goals over the study period. Notably, Goal 9 (Industry, innovation, and infrastructure) had the lowest observed average score (29.7 in the year 2000), but it demonstrated a consistent increase over time, reaching an average of 55.3 such that, in 2022, all average scores were above 50. Conversely, goals 12, 13, and 16 (responsible production and consumption; climate action; and peace, justice, and strong institutions) did not exhibit significant changes over the years. This result may be attributed to socio-economic circumstances.
@@ 3.2 Principal Component analysis
We employed both methods in tandem to analyze SDG data from 2000 to 2022. Specifically, we provided t-SNE with the 10 principal components derived from PCA, which together account for 95% of the variance, we use 50 as perplexity input (this term is explained in the appendix D.0.1) . Subsequently, we applied the DBSCAN algorithm to identify clusters within the reduced t-SNE space. This approach yielded more distinct and cohesive clusters among regions with similar SDG patterns, while also highlighting the separations between regions with different patterns (Figure 4).
@@
A detailed t-SNE plot colored by clusters determined using the DBSCAN algorithm. We use transparency for most dots to reduce visual clutter, highlighting only those that change between clusters with full color. These changes are further emphasized by connecting the highlighted dots with dotted lines. The countries in each clusters are enumerated in Table B.1 in appendix B.
@@
Countries grouped within each cluster are presented in Table B.1 (in Appendix B). Interestingly, the clusters are primarily organized by geographical proximity (see Figure 5), though other influencing factors may include the development level of the countries and existing commercial agreements between them. Cluster 0 predominantly consists of countries from Asia and Africa, with some representation from the Americas, Europe, and Oceania. In contrast, Clusters 1, 2, and 3 are more homogeneous, comprising African, American, and European countries, respectively. Cluster 4 includes a mix of African and Asian countries, while Cluster 5, the smallest, consists of seven developed countries.
@@
Colored map from countries in Table B.1. Cluster 0 is in red, cluster 1 brown, cluster 2 green, cluster 3 purple, cluster 4 light red, cluster 5 blue.
@@
Clustering allowed us to identify distinct regional patterns concerning the United Nations SDGs from 2000 to 2022 (see Figure 6). In this figure standardized SDG scores are presented (as described in the Methods, PCA section), where a value of zero on the y-axis corresponds to the mean SDG value across the dataset. Scores above zero indicate performance above the average for a given SDG, while scores below zero reflect performance below the average. Cluster zero is particularly heterogeneous, encompassing countries from all continents with diverse SDG patterns (as shown by the variability of the gray lines in Figure 6), resulting in average standardized SDG scores that fluctuate around the mean. Other clusters display greater homogeneity. In Cluster 1, the annual averages for most SDGs are negative, with the exceptions of SDG 12 (responsible production and consumption), SDG 13 (climate action), SDG 14 (life below water), and SDG 15 (life on land). Cluster two shows predominantly positive annual averages for the SDGs, except for SDG 10 (reduced inequalities), SDG 15, and SDG 16 (peace, justice, and strong institutions). In Cluster 3, all SDG annual averages exceed the mean, except for SDG 12 and SDG 13, which contrasts with Cluster 4, where the annual scores are below the mean for all SDGs except SDGs 12 and 13. Cluster 5 is similar to Cluster three, but in addition to SDGs 12 and 13, SDGs 14 and 15 are also below the average.
@@ 3.4 Analysis of correlations between SDG
If two goals are correlated positively, it might be that one facilitates the other, or similar conditions promote the change in both of them. If there is a negative correlation, then one hinders the other, or it might be that the conditions that favor one limit the other one. Goals would have independent conditions if their correlation is close to zero.
@@
This observation is significant. It suggests that measures and behavioral changes aimed at reducing superfluous production and consumption could contribute to mitigating climate change but may simultaneously hinder progress towards the other SDGs.
This finding underscores the complexity of sustainability as a system, where the elements interact in a nonlinear fashion, and the overall behavior cannot be understood as a simple summation of individual effects.

### v2

@@ Abstract
The 2030 Agenda for Sustainable Development of the United Nations outlines 17 goals for countries of the world to address global challenges in their development. However, the progress of countries towards these goal has been slower than expected and, consequently, there is a need to investigate the reasons behind this fact. In this study, we have used a novel data-driven methodology to analyze time-series data for over 20 years (2000-2022) from 107 countries using unsupervised machine learning (ML) techniques. Our analysis reveals strong positive and negative correlations between certain SDGs (Sustainable Development Goals). Our findings show that progress toward the SDGs is heavily influenced by geographical, cultural and socioeconomic factors, with no country on track to achieve all the goals by 2030. This highlights the need for a region-specific, systemic approach to sustainable development that acknowledges the complex interdependencies between the goals and the variable capacities of countries to reach them. For this our machine learning based approach provides a robust framework for developing efficient and data-informed strategies to promote cooperative and targeted initiatives for sustainable progress.
@@ 1 Introduction
In 2015, the 2030 Agenda for Sustainable Development Goals (SDG) of the United Nations [1] was established for countries of the world to follow 17 SDGs, 169 targets and 231 globally comparable indicators in their development process (for details, see the Supplementary material S1 section). Since then, the world has faced several setbacks, such as the COVID-19 pandemic, acceleration of climate change, and serious armed conflicts that have slowed down this process. For these reasons and because the SDGs are variably interlinked or interrelated with each other, it seems difficult or even unlikely for countries to achieve the SDGs by 2030. Furthermore, in 2023 an independent group of scientists appointed by the UN to assess the progress of the world in achieving the SDGs and to recommend how to move forward, stated [2]: “The world is not on track to achieve any of the 17 SDGs and cannot rely on change to happen organically. Without accelerated action, the ambitious plan that the world signed up to in 2015 will fail.”
@@
Therefore, it seems obvious that we lack understanding of the complex inter-linkages between different SDGs and what are the drivers leading to their changes. This in turn points out the need for more research, which is made possible by the vast amount of data that the 231 global indicators of the 17 SDGs have produced so far. A thorough analysis of these data is likely to give us some insight into the correlations between, e.g., armed conflicts, climate change, social inequalities, global crime, etc. This would serve as a necessary precursor for the next steps towards sustainability.
@@
As direct interpretation of the available raw data and because the roles and interlinkages between SDGs are difficult to decipher, our aim in this study is to get deeper understanding and insight into them using a modern data analysis approach. We will do this by using unsupervised machine learning methods as they can reveal groups of countries with similar behavior as regards to the 17 SDG’s scores. For example, in the case of diversity, even subtle differences between countries otherwise quite similar should be considered when proposing a new policy agenda. All this should provide valuable information to decision makers at the national and local levels to make well-informed policy choices for achieving the goals and aligning with the 2030 Agenda. The best practices for decision making involve identifying challenges, using data to form insight, and to employ risk analysis while taking into account environmental constraints and local conditions [3].
@@
Sustainability, like many other human endeavors, can be represented as a dynamic and complex network with nonlinear interactions and interdependencies between the Sustainable Development Goals [4]. Therefore, as the SDGs are interconnected and influenced by the intricate web of environmental, political, economic, cultural and social factors, they should not be considered in isolation. Recent progress in complex systems research has enhanced our ability to analyze sustainability using new systemic approaches [5]. Here we will do this by employing the network science approach and its tools, while recognizing that cultural factors play an important role and are difficult to isolate within the global indicators.
@@
So far several studies with different degree of coverage and thoroughness have been conducted to analyze the interlinkages between the SDGs [6, 7, 8, 9, 10]. Among them Ref. [6] constitutes a comprehensive review of 70 previous papers, however, as it is dated for 2020 it does not cover the effect of such recent events as COVID-19. The focus in ref. [7] is very local using only Mexican data. In a number of other studies various methodological approaches have been used, e.g., the nonparametric Spearman’s rank correlation analysis for assessing monotonic relationships between all possible combinations of the indicator data pairs [8], the Multiple Factor Analysis (MFA) used to measure the correlation between indicators [9], and the network analysis approach [10]. In contrast to these methods we will in our study use a combination of techniques to study in detail these interlinkages, which can be either positive or negative, indicating that progress on one goal can enhance or hamper the progress on other goals. For a country, identifying these synergies and trade-offs before setting its policy priorities is crucial to avoid creating new problems while solving existing ones. This interesting issue of SDG interlinkings has been examined in a recent review study [11], which concluded that our understanding of sustainability is still quite rudimentary.
@@
In our study, we will use an unsupervised machine learning multimethod approach with three objectives in mind: i) to examine the changes over time in the 17 SDGs for all countries, identifying possible classifications and discovering similarities and differences in their patterns; ii) to understand how the SDGs correlate across different countries; and iii) to investigate the collective dynamics of similar countries to measure the rate of progress towards achieving the SDGs has been generated in the 21stcentury. Our methods uncover distinct SDG patterns associated with well-defined clusters of countries, each showing different dynamical behaviors towards ideal scores.
@@ 2 Materials and methods
In order to explore, analyse, and understand the nature of interdependences in the complex dataset of SDGs we have devised a novel three-stage unsupervised learning pipeline, depicted in Fig. 1. It consist of dimensionality reduction and clustering by which we are able to effectively simplify and analyze the high-dimensional data, thus making hidden patterns more evident. The use of Principal Component Analysis (PCA)  [12, 13] helps us to reduce the complexity of the data by identifying the main trends and providing a global overview, while t-Distributed Stochastic Neighbor Embedding (t-SNE) [14, 15] focuses on preserving local relationships, thus allowing for a more detailed exploration of similarities between countries. Finally, the DBSCAN clustering technique [16] helps to identify distinct groups, without prior assumptions about the number or shape of clusters, making it suitable for real data, as well as for artificial data. We believe that this combination of analysis methods is the most suitable in providing a global insight into the information embedded in the complex datasets.
Fig 1: A three stage unsupervised learning pipeline for data analysis. The first stage involves data preprocessing, where the dataset undergoes standardization, normalization, and cleaning procedures to address missing values and noise. The second stage consists of sequential dimensionality reduction by PCA to identify principal components to capture global data structure, followed by t-SNE which preserves local relationships in the reduced dimensional space. The final stage applies a clustering algorithm (DBSCAN) to identify distinct groups within the processed data. Countries in each cluster are mapped and the mean trajectories towards ideal scores are calculated. Arrows between the stages indicate the sequential flow of data through the pipeline, with the output of each stage serving as the input for the subsequent stage.
@@ 2.1 Data
The data in this study was collected from the “Online database for the Sustainable Development Report 2023”[17], including only countries without missing data. The original data include 166 countries, but 15 of them lack information on Goal 1 (no poverty), 17 countries did not have information about Goal 10 (reduce inequalities), and 40 countries that lack coastal regions do not have scores for Goal 14 (life below water). Therefore, only 107 countries were included in our analysis.
In this study we used also data of Gross Domestic Product per capita (GDP/c) of every country for the 2022 collected from “Worldometer” [18]. This data serves as a measure of each country’s economic activity as well as overall capability to achieve SDGs.
2.2 Principal Component Analysis (PCA).
@@
PCA was performed using standardized data, that is, the mean and standard deviations were calculated for each SDG throughout the years 2000 to 2022. The corresponding mean was then subtracted from each observation, and the result was divided by the SDG’s standard deviation.
2.3 t-Distributed Stochastic Neighbor Embedding (t-SNE) and clustering.
The t-SNE [14, 15] algorithm is a powerful technique for visualizing high-dimensional data by mapping it into a lower-dimensional space, typically into 2 or 3 dimensions. The primary objective of t-SNE is to represent the underlying structure of the data by placing similar data points close together and dissimilar points further apart in the lower-dimensional space. This is achieved by converting the pairwise Euclidean distances between data points into conditional probabilities, which reflect the likelihood of points being neighbors in the original high-dimensional space. The resulting visualization preserves the local relationships within the data, making it easier to identify patterns and clusters. A more in depth explanation of the method can be found in Supplementary material S1 section.
@@
In our analysis, we applied the t-SNE dimensionality reduction technique followed by the DBSCAN clustering algorithm, carefully tuning their key parameters to ensure optimal results. For t-SNE, the primary parameter is perplexity, which influences the number of effective neighbors for each point and balances the local and global aspects of the data. We selected a perplexity value of 50 after conducting a grid search across a range of values (5 to 50). This value turned out to provide the best representation for our dataset by effectively capturing local relationships between countries while preserving broader trends in their Sustainable Development Goal (SDG) performance. The chosen perplexity was visually validated by examining the resulting 2D t-SNE plots, where the clusters were well separated and meaningful. We also experimented with higher-dimensional t-SNE embeddings and found that similar clustering patterns emerged, indicating the robustness of our approach. Once the t-SNE dimensionality reduction was performed, we used DBSCAN for clustering directly in the 2D t-SNE plane. The main parameter in DBSCAN, epsilon, defines the neighborhood radius within which points are considered part of the same cluster. We tested various values for epsilon and selected the one that resulted in well-defined clusters with minimal noise points, ensuring that the clusters captured meaningful patterns in the data. The chosen epsilon parameter was correlated visually with the 2D t-SNE plots to identify distinct groups of countries based on their SDG trajectories. Similar clustering results were also observed when using higher-dimensional t-SNE projections, confirming the stability of our clustering approach across different dimensionalities. By carefully tuning these parameters, we achieved a balance between local structure and global separations, resulting in insightful and robust clusters that reveal distinct SDG progress patterns among countries.
@@ 3.1 Global view of SDG scores
Given that the effectiveness of sustainability relies on global collaboration, our initial findings focus on the average Sustainable Development Goals (SDGs) scores at a global scale. Figure  2 illustrates the yearly mean scores for each goal, enabling simultaneous observation of the values of all 17 goals over the study period. Notably, Goal 9 (Industry, innovation, and infrastructure) had the lowest observed average score (29.7 in the year 2000), but it demonstrated a consistent increase over time, reaching an average of 55.3 such that, in 2022, all average scores were above 50. Conversely, goals 12, 13, and 16 (responsible production and consumption; climate action; and peace, justice, and strong institutions) did not exhibit significant changes over the years. This result may be attributed to socio-economic circumstances.
@@ 3.2 Principal Component analysis
We employed both methods in tandem to analyze SDG data from 2000 to 2022. Specifically, we provided t-SNE with the 10 principal components derived from PCA, which together account for 95% of the variance, we use 50 as perplexity input (this term is explained in Supplementary material S1 section) . Subsequently, we applied the DBSCAN algorithm to identify clusters within the reduced t-SNE space. This approach yielded more distinct and cohesive clusters among regions with similar SDG patterns, while also highlighting the separations between regions with different patterns (Figure 5).
@@
A detailed t-SNE plot colored by clusters determined using the DBSCAN algorithm. We use transparency for most dots to reduce visual clutter, highlighting only those that change between clusters with full color. These changes are further emphasized by connecting the highlighted dots with dotted lines. The countries in each clusters are enumerated in S1 Table 2.1 (in Supplementary material S1 section)
@@
Countries grouped within each cluster are presented in S1 Table 2.1 (in Supplementary material S1 section). Interestingly, the clusters are primarily organized by geographical proximity (see Figure 6), though other influencing factors may include the development level of the countries and existing commercial agreements between them. Cluster 0 predominantly consists of countries from Asia and Africa, with some representation from the Americas, Europe, and Oceania. In contrast, Clusters 1, 2, and 3 are more homogeneous, comprising African, American, and European countries, respectively. Cluster 4 includes a mix of African and Asian countries, while Cluster 5, the smallest, consists of seven developed countries.
@@
Colored map from countries in S1 Table 2.1 (in Supplementary material S1 section). Cluster 0 is in red, cluster 1 brown, cluster 2 green, cluster 3 purple, cluster 4 light red, cluster 5 blue.
@@
Clustering allowed us to identify distinct regional patterns concerning the United Nations SDGs from 2000 to 2022 (see Figure 7). In this figure standardized SDG scores are presented (as described in the Methods, PCA section), where a value of zero on the y-axis corresponds to the mean SDG value across the dataset. Scores above zero indicate performance above the average for a given SDG, while scores below zero reflect performance below the average. Cluster zero is particularly heterogeneous, encompassing countries from all continents with diverse SDG patterns (as shown by the variability of the gray lines in Figure 7), resulting in average standardized SDG scores that fluctuate around the mean. Other clusters display greater homogeneity. In Cluster 1, the annual averages for most SDGs are negative, with the exceptions of SDG 12 (responsible production and consumption), SDG 13 (climate action), SDG 14 (life below water), and SDG 15 (life on land). Cluster two shows predominantly positive annual averages for the SDGs, except for SDG 10 (reduced inequalities), SDG 15, and SDG 16 (peace, justice, and strong institutions). In Cluster 3, all SDG annual averages exceed the mean, except for SDG 12 and SDG 13, which contrasts with Cluster 4, where the annual scores are below the mean for all SDGs except SDGs 12 and 13. Cluster 5 is similar to Cluster 3, but in addition to SDGs 12 and 13, SDGs 14 and 15 are also below the average.
We note that the Clusters differ not only in geographical localization and SDG patterns but also in their average GDP/c (Gross Domestic Product per capita), a widely used economic indicator. Among the included countries, those belonging to Cluster 1 have the lowest average GDP/c (\$1,$4$$\pm$\$883.72). The average GDP/c value for the countries in Cluster 4 is \$3,941.89 $\pm$ \$3,113.04 while for the countries grouped in Cluster 0 it is \$7,111.55 $\pm$ \$4,150.60, and similar to the one observed for Cluster 2 (\$10,328.50 $\pm$ \$4,923.47). Clusters 3 and 5 have statistically equal average CDP/c values, i.e., \$44,846.74 $\pm$ \$25,956.59 and \$43,524.00 $\pm$ \$19,122.98 respectively.
@@ 3.4 Analysis of correlations between SDG
If two goals are positively correlated, it might be that one facilitates the other or similar conditions promote the change in both of them. If there is a negative correlation, then one hinders the other, or it might be that the conditions that favor one limit the other one. The goals would have independent conditions if their correlation is close to zero.
@@
This observation is significant. It suggests that measures and behavioral changes aimed at reducing superfluous production and consumption could contribute to mitigating climate change but may simultaneously hinder progress towards the other SDGs. The fact that these goals are anti-correlated with most other goals could depend on various economic factors like how the production is structured to use the resources of the planet. To gain a deeper understanding of this, we would need more detailed data on the economic structure of countries; therefore, this is beyond the present study based on the SDGs only. This finding underscores the complexity of sustainability as a system, where the elements interact in a nonlinear fashion, and the overall behavior cannot be understood as a simple summation of individual effects.